OpenStudy (anonymous):

A rectangular paperboard 28in long and 20in wide has a semicircle cut out of it, as shown below. Find the area of the paperboard that remains. Use the value 3.14 for π , and do not round your answer. Be sure to include the correct unit in your answer.

OpenStudy (anonymous):

I know I need to find the area of the semi circle first to subtract from the over all area and find the remaining piece but how do I go about that again? =o

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

the area of the entire rectangle = 560 area of semi circle=? area of rectangle - area of semi circle = area of paperboard that remains

OpenStudy (anonymous):

ok so I found the formula for area of a semi circle. A= ( Pi * r^2) / 2 A= (Pi * 10^2) / 2 A = 50Pi

OpenStudy (anonymous):

SOOO, 560 - 50Pi = Ok now this is where I'm stuck because mathway says that's as simplified as it gets, but i'm pretty sure that's not the answer they are looking for and it won't let me type Pi in the answer box =o

OpenStudy (anonymous):

Ok so now I replaced Pi with 3.14 and now I have 560 - 157 = 403 area of remaining board = 403 Is that right?

OpenStudy (anonymous):

you are correct.

OpenStudy (anonymous):

ahh yes! lol thank you =)

OpenStudy (anonymous):

yw